# Hierarchical Object-Oriented POMDP Planning for Object Rearrangement

### v2

@@ Abstract
We present an online planning framework for solving multi-object rearrangement problems in partially observable, multi-room environments. Current object rearrangement solutions, primarily based on Reinforcement Learning or hand-coded planning methods, often lack adaptability to diverse challenges. To address this limitation, we introduce a novel Hierarchical Object-Oriented Partially Observed Markov Decision Process (HOO-POMDP) planning approach. This approach comprises of (a) an object-oriented POMDP planner generating sub-goals, (b) a set of low-level policies for sub-goal achievement, and (c) an abstraction system converting the continuous low-level world into a representation suitable for abstract planning. We evaluate our system on varying numbers of objects, rooms, and problem types in AI2-THOR simulated environments with promising results.
@@ 1 Introduction
Multi-object rearrangement with egocentric vision in realistic home environments is a fundamental challenge in embodied AI, encompassing complex tasks that require perception, planning, navigation, and manipulation. This problem becomes particularly demanding in multi-room settings with partial observability, where large parts of the environment are not visible at any given time. Such scenarios are ubiquitous in everyday life, from tidying up households to organizing groceries, making them critical for the development of next-generation home assistant robots.
@@
Existing approaches to multi-object rearrangement typically fall into two categories: Reinforcement Learning (RL) methods and hand-coded planning systems. RL methods often struggle as the problem becomes increasingly complex and lengthy, making it difficult to scale to more challenging scenarios. To address this limitation, many researchers have adopted a modular approach, decomposing the task into a series of subtasks (Gu et al., 2022) such as manipulation skills, navigation skills, or exploration skills. These subtasks are then sequenced together in different ways to accomplish the overall goal. However, current modular approaches have their own limitations. Some pre-determine the sequence in which to apply skills, while others use greedy planners (Trabucco et al., 2022). This constrains their potential for full optimization in terms of determining the optimal order to interact with objects and handling new problems, such as when another object blocks the path to an object or location or if the goal location itself is obstructed. To overcome these challenges, a more general approach that incorporates high-level planning based on the current state would be beneficial. Such an approach would enable the system to handle novel problems without requiring extensive re-learning from scratch, thus increasing its adaptability and efficiency in diverse environments. Solving these new problems is particularly important in the context of household robots, where obstacles, blocked paths, and obstructed goals are common occurrences that a robust system must be able to handle effectively.
@@
While significant progress has been made in rearrangement, the majority of current research focuses on single-room settings or assumes that a large number of objects are visible at the beginning of the task, either through a third-person bird’s eye view (Ghosh et al., 2022) or a first-person view where most of the room is visible. However, as we move towards the more practical version of the problems, such as cleaning a house, the majority of the space and objects to be manipulated are not initially visible, and existing solutions begin to falter. This scenario of partial observability introduces several major challenges: 1) uncertainty over object locations, as the starting positions of objects are unknown; 2) execution efficiency of searching for objects while simultaneously moving them to the correct goal locations; 3) scalability of planning over increasing numbers of objects and rooms; 4) extensibility to scenarios involving blocked goals or obstructed paths; and 5) graceful handling of object detection failures.
@@
Figure 1: a) shows the agent’s ego-centric view of the world at initialization. b) shows the top-down view of the environment before the start of the rearrangement task. The dotted bounding boxes indicate the starting state of the object, and the solid bounding boxes indicate the object’s goal state. In a) The lines indicate the path between the start and goal states. Object 1’s path to goal is blocked by object 2, and its goal location is blocked by object 3. Object 3’s path to its goal is also blocked by object 2. Hence, object 2 needs to be moved to its goal, then object 3 to its goal, and only then can object 1 be moved to its goal. Objects 4 and 5 are blocking each other’s goals, and hence, one of them needs to be placed elsewhere, and then the swap can happen.
We employ a hierarchical approach to decompose the complex multi-object rearrangement problem into two distinct levels. This structure allows us to address object combinatorics at the high level while managing regional interactions at the low level. By separating these aspects, we effectively mitigate the challenges inherent to each. We use an object-oriented representation rather than low-level perceptual representations such as point clouds, which reduces the complexity of planning and is more natural in our setting. Our hierarchical Object-Oriented Partially Observed Markov Decision Process (HOO-POMDP) planning framework has the following components: 1) A perception module that detects objects and outputs object-oriented observations, 2) A belief update and state abstraction system, 3) A high-level planning module that plans the sequence of locations to be visited or objects to be transported, 4) A low-level navigation and manipulation module that plans a path to its next destination or control actions to pick and place objects.
In more detail, the agent maintains a factored belief state over object locations based on its prior beliefs and the output of its object detection module. An abstract object-oriented POMDP planner produces a high-level policy for the given goal. The first high-level action in the plan is treated as a (navigation or manipulation) sub-goal for the low-level planner, which produces a lower-level sequence of actions. The agent takes the first action in the lower-level plan, updates its belief state based on its new perception, and the cycle repeats.
The main contributions of our research are as follows:
@@
  • A modular planning system, which includes an object-oriented planner and a state abstraction module for object rearrangement in multi-room environments.
@@
  • A new dataset featuring blocked path problems and expanded room configurations alongside existing rearrangement challenges.
  • An empirical evaluation of the system in the new dataset in AI2Thor under different conditions.
@@ 2 Related Work
Rearrangement: Rearrangement is the problem of manipulating the placement of objects by picking, moving, and placing them according to a goal configuration. In this work, we are mainly concerned with the rearrangement of objects by mobile agents in simulated environments such as AI2Thor and Habitat (Kolve et al., 2017),(Szot et al., 2021) and (Gan et al., 2020). We hypothesize that successful rearrangement planning in simulators with noisy sensors and effectors will go a long way towards successful planning and execution in robotics.
@@
There are many versions of the rearrangement problem in literature. In tabletop rearrangement, a robot hand with a fixed base moves objects around to achieve a certain configuration in a limited space (Zhai et al., 2024; Huang et al., 2024). Many current approaches for rearrangement by mobile agents work by finding the misplaced objects and then use greedy planners to decide what order to move the objects in (Gadre et al., 2022; Trabucco et al., 2022; Sarch et al., 2022). This can lead to a high traversal cost since it is not explicitly optimized. The above works and others such as (Mirakhor et al., 2024a) are also limited to a single-room setting where most objects are visible to the agent.
@@
Rearrangement has also been studied from the Task and Motion Planning (TAMP) perspective (Garrett et al., 2020a; 2021; b). perspective. Garrett et al. (2020b) is limited to a single-room kitchen problem and assumes perfect detection of objects. Unlike most previous work, our proposed solution optimizes the traversal cost and addresses multi-room settings and imperfect object detection in an integrated POMDP framework. Tekin et al. (2023) and Mirakhor et al. (2024b) address the multi-room rearrangement problem. However, the decision-making process of Tekin et al. (2023) about when to explore and when to move an object is fixed and assumes perfect object detection. Our framework optimizes, naturally handles exploration and manipulation, and addresses object detector failures in a unified framework. Mirakhor et al. (2024b) make the assumption that objects are always on top of or inside containers. This limits its extendability to handling new problems, such as blocked paths where the objects could be in the path of another object without any containers. Our framework naturally allows for these new possibilities.
Object Oriented POMDP (OOPOMDP) Planning: Our work is partly inspired by Wandzel et al. (2019b), who OOPOMDP and perform a multi-object search in a 2D environment. Zheng et al. (2023) and Zheng et al. (2022) extend this formulation to object search in 3D environments. However, they are limited to the task of object search. In our work, we build on their formulation of OOPOMDP and extend it to include rearrangement actions and their corresponding belief updates. We further extend this rearrangement OO-POMDP to HOO-POMDP through action abstraction.
@@ 3 Problem Formulation
Environment and Agent: Our agent is developed for the AI2Thor simulator environment (Kolve et al., 2017). It consists of a simulated house with a set of objects located in one or more rooms. The agent can take the following low-level actions: $A_{s}$ = ($MoveAhead$, $MoveBack$, $MoveRight$, $MoveLeft$, $RotateLeft$, $RotateRight$, $LookUp$, $LookDown$, $PickObject_{i}$, $PlaceObject$, $Start_{loc},Done)$. The $Move$ actions move the agent by a distance of $0.25m$ in the environment. The $Rotate$ actions rotate the agent pitch by $90$ degrees. The $Look$ actions rotate the agent yaw by $30$ degrees. $Start$ action starts the simulator and places the agent at the given location, and the $Done$ action ends the simulation. After executing any of the above actions, the simulator outputs the following information: a) RGB and Depth images, 2) the agent’s position ($x,y,pitch,yaw$), and 3) whether the action was successful. There are two types of objects in the world - interactable objects and receptacle objects. The interactables are the ones that can be picked and placed. The receptacles are objects that are not movable but can hold other objects.
@@
Task Setup: Rearrangement is done in 2 phases. Walkthrough phase and rearrange phase. The walkthrough phase is meant to get information about stationary objects. The 2D occupancy map is generated in this phase, as well as the corresponding 3D Map. We get the size of the house(width and length) from the environment and uniformly sample points in the environment. We then take steps to reach these locations (if possible - some might be blocked). This simple algorithm ensures we explore the full house. At each of the steps, we receive the RGB and Depth. Using this, we create a 3D point cloud at each step and combine them all to get the overall 3D point cloud of the house with stationary objects. We then discretize this point cloud into 3D map voxels of size 0.25m, we further flatten this 3D map into a 2D map of grid cells (location in the 2D map is occupied if there exists a point at that 2D location at any height in the 3D map). While doing this traversal, we also get information about the receptacles by detector on the RGB images we receive during this traversal. This ends the walkthrough phase. (this walkthrough process needs to be done only once for any house configuration of stationary objects - walls, doors, tables, etc.). Then, objects are placed at random locations (done using AI2Thor environment reinitialization). This is when the rearrangement phase begins, with the planner taking the following as input - the map generated in the walkthrough phase, the set of object classes to move, and their goal locations.
4 Hierarchical Object Oriented POMDP (HOO-POMDP) Planner
@@
This section presents our online planning framework designed to solve multi-object rearrangement problems in partially observable, multi-room simulated home environments. Our approach enables the system to tackle complex rearrangement tasks that involve challenging sub-problems, such as clearing blocked paths.
Figure 2: The agent receives RGB and depth images from environment at the start. The vision module creates the observation from this input and sends it to the belief update system. Belief is updated based on observation, and an abstract state is generated, which is sent to the OO POMDP Planner that outputs sub-goals. The sub-goals are used by the low-level policy executors to get and execute low-level actions in the environment.
@@
Overview: Once the initial list of receptacles and 3D map have been generated, they, along with the goal information, are sent to the HOO-POMDP planner. The system operates in a cyclic fashion, integrating perception, belief update, state abstraction, abstract planning, and action execution (see Figure 2 and Algorithm 1). First, the perception subsystem detects objects in the RGB and depth image and outputs the observation $z$. This $z$ is used by the belief update subsystem to update the object-oriented belief state, which consists of the probability of each object being at a certain location in $M^{2D}$. The abstraction system uses this information to update its abstract state. The updated abstract state is sent to the abstract POMDP planner, which outputs a sub-goal that corresponds to a low-level policy. The low-level policy executor executes the low-level policy corresponding to the sub-goal. This might involve navigating to a specific location, grasping an object, or placing an object in a new position. After each action is executed, the environment state changes. The agent receives new output from the environment, and the cycle repeats until the overall rearrangement task is completed. In the rest of this section, we will discuss each of the subsystems and their interaction.
1 Function HOO_POMDP_Rearrangement():
2       env $\leftarrow$ InitializeEnv (); beliefState $\leftarrow$ InitializeBeliefState ();
3       agent $\leftarrow$ InitializeAgent (); loc $\leftarrow$ random(); lowLevelAction $\leftarrow$ Startloc;
4       while not TaskComplete() do
5             rgb, depth $\leftarrow$ env.GetObservation (lowLevelAction);
6             observation $\leftarrow$ ProcessObservation (rgb, depth);
7             beliefState $\leftarrow$ UpdateBelief (beliefState, observation, lowLevelAction);
8             abstractState $\leftarrow$ GenerateAbstractState (beliefState, G, R);
9             abstractAction $\leftarrow$ POUCTPlanner (abstractState, beliefState);
10             if abstractAction == Done then
11                   return()
12            lowLevelPolicy $\leftarrow$ SelectLowLevelPolicy (abstractAction);
13             lowLevelAction $\leftarrow$ lowLevelPolicy.GetAction (abstractAction, rgb, depth);
14
Algorithm 1 HOO-POMDP Rearrangement($M^{2D},G,R$)
Abstract OO-POMDP Planning :
@@ 1 Function HOO_POMDP_Rearrangement():
Background: POMDP: A POMDP is a 7-tuple $(S,A,T,R,\gamma,O,O_{model})$ (Kaelbling et al., 1998). The state space $S$ is the set of states in which the agent and the objects in the environment can be. Action space $A$ is the set of actions that can be taken in the environment. The transition function $T(s,a,s^{\prime})=p(s^{\prime}|s,a)$ is the probability of reaching the state $s^{\prime}$ when the action $a$ is taken in the current state $s$. The agent receives an observation $z\in O$ when an action is taken. The probability of receiving an observation when being in a given state $s$ after having taken action $a$ is defined by the observation model $O_{model}(s,a,z)=p(z|s,a)$. The reward function $R(s,a)$ defines the reward received when taking action $a$ in state $s$, and $\gamma$ is the discount factor. In a partially observed world, the agent does not know its exact state and maintains a distribution over possible states, i.e., a belief state $b$. The belief is updated when an action is taken, and observation is received with the following equation, where $\eta$ is the normalizing constant:
$b^{\prime}(s^{\prime})=\eta O(s^{\prime},a,z)\sum_{s\in S}T(s,a,s^{\prime})b(s)$ (1)
@@
Object Oriented POMDP: Object-oriented POMDP factors the state and observations over the objects. Each state $s$ is represented as a tuple of its $n$ objects $s=(s_{1},\ldots,s_{n})$ ,each observation $z=(z_{1},\ldots,z_{n})$ (Wandzel et al., 2019a) and the belief state $b$ is factorized as $b=\prod_{i=0}^{n}b_{i}$
Abstract Object Oriented POMDP:
@@
We now instantiate the rearrangement problem as an abstract POMDP. In our definition of the abstract OOPOMDP, we make an object independence assumption - that at any given time, the observation and belief of any object do not depend on any other object.More formally, $(P(z_{i}|s_{j},z_{j},s_{i})=P(z_{i}|s_{i})$, observation $z_{i}$ is independent of the states and observations of other objects, conditioned on its own state $s_{i}$ (observations are conditionally independent). Similarly, we also assume $P(s’_{i}|s_{i},s_{j},a)=P(s’_{i}|s_{i},a)$ when $j!=i$, i.e., the next state of object $i$ only depends on its own previous state and the action. This allows us to represent the state and observation as entities factored on objects, which in turn helps make independent belief updates for each object(Algorithm 2).
  • State Space: We use a factored state space that includes the robot state $s_{r}$, and the target object states $s_{targets}$. The complete state is represented as $s=(s_{r},\ s_{targets})$. $s_{targets}=$ ($s_{target_{1}},\ldots,s_{target_{n}})$ where $n$ is the number of objects to be moved. $s_{target_{i}}=(loc_{i}$$pick_{i}$$place_{locs}$$is\_held$$at\_goal$$g_{i}):loc_{i}$ is the current location of the object, $pick_{i}$ corresponds to the location from where this object can be picked, $place_{locs}$ corresponds to the set of locations (absolute 2D coordinates) from where this object can be placed from. $g_{i}$ is the goal location of the object. All locations are discretized grid coordinates in the 2D map ($M^{2D}$).
@@
    • $\textit{Move}_{AB}$ The move action moves the agent from location A to location B.
    • $\textit{Rotate}_{angle}$ The rotate action rotates the agent to a given angle.
    • $\textit{PickPlace}_{Object_{i}-goal_{loc}}$ The PickPlace action picks $Object_{i}$ from the current position of the robot and places it at the given $goal_{loc}$.
  • Observation Space: We use a factored observation space similar to state space factorization. Each observation can be divided into the robot observation and object observation $z=(z_{robot},\ z_{objects})$, where $z_{objects}=(z_{target_{1}},\ \ldots\ z_{target_{n}})$. Each observation $z_{target_{i}}\in L\cup\text{Null}$ is a detection of the object $i$’s location or Null based on the detector’s output for object $i$.
@@
    $=\prod_{i=1}^{n}\Pr(z_{target_{i}}|s_{target_{1}},\ldots,s_{target_{n}},s_{\text{r}}),(\text{all }z_{target_{i}}\text{ are independent })$ (3)
    $=\prod_{i=1}^{n}\Pr(z_{target_{i}}|s_{target_{i}},s_{\text{r}}),(z_{target_{i}}\text{ does not depend on {state} of other objects})$ (4)
@@
    • $\textit{Move}_{AB}$ : The cost of moving from location A to B [$Cost=-1*N_{a}$ (where $N_{a}$ number of required actions)].

### v3

@@ Abstract
We present an online planning framework and a new benchmark dataset for solving multi-object rearrangement problems in partially observable, multi-room environments. Current object rearrangement solutions, primarily based on Reinforcement Learning or hand-coded planning methods, often lack adaptability to diverse challenges. To address this limitation, we introduce a novel Hierarchical Object-Oriented Partially Observed Markov Decision Process (HOO-POMDP) planning approach. This approach comprises of (a) an object-oriented POMDP planner generating sub-goals, (b) a set of low-level policies for sub-goal achievement, and (c) an abstraction system converting the continuous low-level world into a representation suitable for abstract planning. To enable rigorous evaluation of rearrangement challenges, we introduce MultiRoomR, a comprehensive benchmark featuring diverse multi-room environments with varying degrees of partial observability (10-30% initial visibility), blocked paths, obstructed goals, and multiple objects (10-20) distributed across 2-4 rooms. Experiments demonstrate that our system effectively handles these complex scenarios while maintaining robust performance even with imperfect perception, achieving promising results across both existing benchmarks and our new MultiRoomR dataset.
@@ 1 Introduction
Multi-object rearrangement with egocentric vision in realistic simulated home environments is a fundamental challenge in embodied AI, encompassing complex tasks that require perception, planning, navigation, and manipulation. This problem becomes particularly demanding in multi-room settings with partial observability, where large parts of the environment are not visible at any given time. Such scenarios are ubiquitous in everyday life, from tidying up households to organizing groceries, making them critical for the development of next-generation home assistant robots.
@@
Existing approaches to multi-object rearrangement typically fall into two categories: Reinforcement Learning (RL) methods and hand-coded planning systems. RL methods (Weihs et al., 2021) struggle with scaling to complex scenarios, while modular approaches that decompose tasks into subtasks (Gu et al., 2022) have different limitations. Some use predetermined skill sequences, while others employ greedy planners (Trabucco et al., 2022), restricting their ability to determine optimal object interaction orders or to handle novel problems such as blocked paths and obstructed goals. A more general approach that incorporates high-level planning would enable systems to handle these challenges without extensive retraining, particularly important for household robots operating in environments where such obstacles are common.
@@
Although significant progress has been made in rearrangement, the majority of current research focuses on single-room settings or assumes that a large number of objects are visible at the beginning of the task, either through a third-person bird’s eye view (Ghosh et al., 2022) or a first-person view where most of the room is visible (Trabucco et al., 2022). However, as we move towards the more practical version of the problems, such as cleaning a house, the majority of the objects to be manipulated are not initially visible, and existing solutions begin to falter. Rearrangement in realistic multi-room environments introduces several key challenges: 1) uncertainty over object locations, as the initial positions of objects are unknown; 2) execution efficiency of searching for objects while simultaneously moving them to the correct goal locations; 3) scalability of planning over increasing numbers of objects and rooms; 4) extensibility to scenarios involving blocked goals or blocked paths; and 5) graceful handling of object detection failures. To enable rigorous evaluation of these challenges, we introduce MultiRoomR, a comprehensive benchmark dataset featuring diverse multi-room environments with varying degrees of partial observability, blocked paths, obstructed goals, and multiple objects distributed across rooms—scenarios that existing datasets like RoomR (Weihs et al., 2021) do not adequately represent.
@@
Figure 1: a) Agent’s ego-centric view at initialization. b) Top-down view of the environment showing object starting positions (dotted boxes) and goal positions (solid boxes). Lines indicate paths between start and goal states. Object 1’s path is blocked by object 2, and its goal location is blocked by object 3. Object 3’s path is also blocked by object 2, requiring a specific sequence: move object 2 first, then object 3, and finally object 1. Objects 4 and 5 block each other’s goals, requiring one to be temporarily placed elsewhere before completing the swap.
To effectively tackle these complex challenges in multi-room settings, we propose the Hierarchical Object-Oriented Partially Observable Markov Decision Process (HOO-POMDP) framework, a modular approach to the multi-object rearrangement problem. This framework consists of two key components: a high-level POMDP planner that reasons under uncertainty and a set of specialized low-level policies for executing specific tasks.
The high-level planning with uncertainty enables joint optimization over exploration and object manipulation decisions, determining when to search for unseen objects and when to rearrange detected ones based on the current belief state. Meanwhile, the low-level policies handle the execution details of navigation and manipulation, freeing the high-level planner from concerns about continuous action spaces and perceptual representations. This separation of concerns allows each component to focus on its strengths—strategic decision-making at the high level and specialized task execution at the low level. Our main contributions include:
@@
  • A modular planning system, which includes an object-oriented POMDP planner and a state abstraction module for object rearrangement in multi-room environments.
@@
  • A new dataset MultiRoomR featuring blocked path problems and expanded room configurations alongside existing rearrangement challenges.
  • An empirical evaluation of the system in an existing and the new MultiRoomR dataset in AI2Thor.
@@ 2 Related Work
Rearrangement: Rearrangement is the problem of manipulating the placement of objects by picking, moving, and placing them according to a goal configuration. In this work, we are mainly concerned with the rearrangement of objects by mobile agents in simulated environments such as AI2Thor (Kolve et al., 2017) and Habitat (Szot et al., 2021) and ThreeDWorld (Gan et al., 2020). There are many versions of the rearrangement problem in literature. In tabletop rearrangement, a robot hand with a fixed base moves objects around to achieve a certain configuration in a limited space (Zhai et al., 2024; Huang et al., 2024). Many current approaches for rearrangement by mobile agents work by finding the misplaced objects and then use greedy planners to decide what order to move the objects in (Gadre et al., 2022; Trabucco et al., 2022; Sarch et al., 2022). This can lead to a high traversal cost since it is not explicitly optimized. The above works and others such as (Mirakhor et al., 2024a) are also limited to a single-room setting where most objects are visible to the agent.
@@
Rearrangement has also been studied from the Task and Motion Planning (TAMP) perspective (Garrett et al., 2020a, 2021, b). Garrett et al. (2020b) is limited to a single-room kitchen problem and assumes perfect detection of objects. Unlike most previous work, our proposed solution optimizes the traversal cost and addresses multi-room settings and imperfect object detection in an integrated POMDP framework. Tekin et al. (2023) and Mirakhor et al. (2024b) address the multi-room rearrangement problem. However, the decision process of when to explore and when to move an object in (Tekin et al., 2023) is fixed and assumes perfect object detection. Our unified framework optimally combines exploration and manipulation, and naturally addresses object detector failures. Mirakhor et al. (2024b) assume that objects are always on top of or inside containers. This limits its extendability to handling new problems, such as blocked paths where the objects could be in the path of other objects and outside containers. Our framework naturally allows for these possibilities. Large language models (LLMs) have also been used to solve rearrangement (Chang et al., 2024), but the advantage of a planner is that it provides a completeness guarantee - given enough time, the planner will find a solution whereas an LLM does not provide the same guarantee.
POMDP Planning: Our work builds upon Wandzel et al. (2019)’s object-oriented POMDP (OO-POMDP) for 2D multi-object search, later extended to 3D environments by Zheng et al. (2023) and Zheng et al. (2022). While these OO-POMDP approaches are limited to search tasks, we extend the formulation to include rearrangement actions with corresponding belief updates. Our HOO-POMDP further introduces action abstraction, distinguishing it from existing hierarchical POMDP work (Serrano et al., 2021) which lacks object-oriented belief maintenance. This combination of hierarchical planning with object-oriented beliefs enables efficient planning for complex rearrangement tasks that would be intractable in flat POMDP representations.
@@ 3 Problem Formulation
Environment and Agent: Our agent is developed for the AI2Thor simulator environment (Kolve et al., 2017). It consists of a simulated house with a set of objects located in one or more rooms. The agent can take the following low-level actions: $A_{s}$ = ($MoveAhead$, MoveBackMoveBack, MoveRightMoveRight, MoveLeftMoveLeft, RotateLeftRotateLeft, RotateRightRotateRight, LookUpLookUp, LookDownLookDown, $PickObject_{i}$, $PlaceObject$, $Start_{loc},Done)$. The $Move$ actions move the agent by a distance of $0.25m$. The $Rotate$ actions rotate the agent pitch by $90$ degrees. The $Look$ actions rotate the agent yaw by $30$ degrees. $Start$ action starts the simulator and places the agent at the given location, and the $Done$ action ends the simulation. After executing any of the actions, the simulator outputs the following information: a) RGB and Depth images, 2) the agent’s position ($x,y,pitch,yaw$), and 3) whether the action was successful. There are two types of objects in the world - interactable objects that can be picked and placed, and receptacle objects that are not movable but can hold other objects.
@@
Task Setup: Rearrangement is done in 2 phases. Walkthrough phase and rearrange phase. The walkthrough phase is meant to get information about stationary objects. The 2D occupancy map is generated in this phase, as well as the corresponding 3D Map. We get the size of the house (width and length) from the environment and uniformly sample points in the environment. We then take steps to reach these locations (if possible - some might be blocked). This simple algorithm ensures we explore the full house. At each of the steps, we receive the RGB and Depth. Using this, we create a 3D point cloud at each step and combine them all to get the overall 3D point cloud of the house with stationary objects. We then discretize this point cloud into 3D map ($M^{3D}$) voxels of size 0.25m, we further flatten this 3D map into a 2D map ($M^{2D}$) of grid cells (location in the 2D map is occupied if there exists a point at that 2D location at any height in the 3D map). While doing this traversal, we also get information about the receptacles by detector on the RGB images we receive during this traversal. This ends the walkthrough phase, which needs to be done only once for any house configuration of stationary objects - walls, doors, tables, etc. Then, objects are placed at random locations (done using AI2Thor environment reinitialization). This is when the rearrangement phase begins, with the planner taking the following as input: the map generated in the walkthrough phase, the set of object classes to move, and their goal locations.
4 Hierarchical Object Oriented POMDP (HOO-POMDP) Framework
@@
This section presents our online planning framework designed to solve multi-object rearrangement problems in partially observable, multi-room simulated home environments.
@@
Overview: Once the initial list of receptacles and $M^{3D}$ have been generated, they, along with the goal information, are sent to the HOO-POMDP planner. The system operates in a cyclic fashion, integrating perception, belief update, state abstraction, abstract planning, and action execution (see Figure 2 and Algorithm 1). First, the perception system detects objects in the RGB and depth image and outputs the observation $z$, which is used by the belief update system to update its belief state. The belief state consists of the probability of each object being at a certain location in $M^{2D}$. The abstraction system uses this information to update its abstract state. The updated abstract state is sent to the abstract POMDP planner, which outputs a sub-goal that corresponds to a low-level policy. The low-level policy executor executes the low-level policy corresponding to the sub-goal. This might involve navigating to a specific location, grasping an object, or placing an object in a new position. After each action is executed, the environment state changes. The agent receives new output from the environment, and the cycle repeats until the overall rearrangement task is completed. In the rest of this section, we will discuss each of the subsystems and their interaction.
Figure 2: The agent receives RGB and depth images from environment at the start. The vision module creates the observation from this input and sends it to the belief update system. Belief is updated based on observation, and an abstract state is generated, which is sent to the $O$ POMDP Planner that outputs sub-goals. The sub-goals are used by the low-level policy executors to plan and execute low-level actions in the environment.
@@
Background: POMDP: A POMDP is a 7-tuple $(S,A,T,R,\gamma,O,O_{model})$ (Kaelbling et al., 1998). The state space $S$ is the set of states in which the agent and the objects in the environment can be. Action space $A$ is the set of actions that can be taken in the environment. The transition function $T(s,a,s^{\prime})=p(s^{\prime}|s,a)$ is the probability of reaching the state $s^{\prime}$ when the action $a$ is taken in the current state $s$. The probability of observing $z\in O$ after having taken action $a$ in a state $s$ is defined by the observation model $O_{model}(s,a,z)=p(z|s,a)$. The reward function $R(s,a)$ defines the reward received when taking action $a$ in state $s$, and $\gamma$ is the discount factor. In a partially observed world, the agent does not know its exact state and maintains a distribution over possible states, i.e., a belief state $b$. The belief is updated when an action $a$ is taken, and observation $z$ is received with the following equation, where $\eta$ is the normalizing constant:
$b^{\prime}(s^{\prime})=\eta O_{model}(s^{\prime},a,z)\sum_{s\in S}T(s,a,s^{\prime})b(s)$ (1)
@@
Object Oriented POMDP: Object-oriented POMDP factors the state and observations over the objects. Each state $s$ is represented as a tuple of its $n$ objects $s=(s_{1},\ldots,s_{n})$, each observation $z=(z_{1},\ldots,z_{n})$ and the belief state $b$ is factorized as $b=\prod_{i=0}^{n}b_{i}$ (Wandzel et al., 2019).
Abstract Object Oriented POMDP formulation:
@@
We now instantiate the rearrangement problem as an abstract POMDP. In our definition of the abstract OOPOMDP, we make an object independence assumption - that at any given time, the observation and state of any object do not depend on any other object. More formally, $P(z_{i}|s_{j},z_{j},s_{i})=P(z_{i}|s_{i})$, observation $z_{i}$ is independent of the states and observations of other objects, conditioned on its own state $s_{i}$. Similarly, we also assume $P(s_{i}^{\prime}|s_{i},s_{j},a)=P(s_{i}^{\prime}|s_{i},a)$ where $j\neq i$, i.e., the next state of object $i$ only depends on its own previous state and the action. This allows us to represent the state and observation factored by objects, which in turn helps make independent belief updates for each object (Algorithm 1).
[…]
  1. 1.
    $env\leftarrow$ InitializeEnv() ;
  2. 2.
    $agent\leftarrow$ InitializeAgent()
  3. 3.
    $belief\leftarrow$ InitializeBeliefState()
  4. 4.
    $loc\leftarrow$ random()
  5. 5.
    $lowLevelAct\leftarrow$ Startloc
  6. 6.
    while not TaskComplete() do
  7. 7.
    $rgb,depth\leftarrow$ envenv.Execute(lowLevelActlowLevelAct)
  8. 8.
    $obs\leftarrow$ PerceptionSystem($rgb,depth$)
  9. 9.
    $belief\leftarrow$ BeliefUpdate($belief,obs,$
    lowLevelActlowLevelAct)
  10. 10.
    $absState\leftarrow$ GenerateAbsState($belief,$
    $G,R$)
  11. 11.
    $absAct\leftarrow$ POUCTPlanner($absState,$
    beliefbelief)
  12. 12.
    if $absAct=$ Done then
  13. 13.
    return
  14. 14.
    $lowLevelPolicy\leftarrow$ GetLowLevelPol-
    icy(absActabsAct)
  15. 15.
    $lowLevelAct\leftarrow$ lowLevelPolicylowLevelPolicy.GetAct
      ($absAct,rgb,depth$)
Algorithm 1 HOO-POMDP framework
  1. 1.
    Input: beliefState $b$, observation $z$, action $a$
  2. 2.
    for each object $i$ in $b$ do
  3. 3.
    for each possible state $s_{ij}$ of object $i$ do
  4. 4.
    if action is navigation then
  5. 5.
    $b^{\prime}_{i}(s_{ij})\leftarrow p(z_{i}|s_{ij})b_{i}(s_{ij})$
  6. 6.
    else if action is place then
  7. 7.
    $b^{\prime}_{i}(s_{ij})\leftarrow\begin{cases}1&\text{if }s_{ij}=\text{action.goalLocation}\\ 0&\text{otherwise}\end{cases}$
  8. 8.
    else if action is pick then
  9. 9.
    $b^{\prime}_{i}(s_{ij})\leftarrow\begin{cases}1&\text{if }s_{i}=\text{action.agentLocation}\\ 0&\text{otherwise}\end{cases}$
  10. 10.
    end if
  11. 11.
    end for
  12. 12.
    end for
  13. 13.
    return $b^{\prime}$
Algorithm 2 BeliefUpdate
@@
    • $\textit{Move}_{AB}$ - The move action moves the agent from location A to location B.
    • $\textit{Rotate}_{angle}$ - The rotate action rotates the agent to a given angle.
    • $\textit{PickPlace}_{Object_{i}-goal_{loc}}$ - The PickPlace action picks $Object_{i}$ from the current position of the robot and places it at the given $goal_{loc}$.
  • Observation Space: We use a factored observation space similar to state space factorization. Each observation can be divided into the robot observation and object observation $z=(z_{robot},\ z_{objects})$, where $z_{objects}=(z_{target_{1}},\ \ldots\ z_{target_{n}})$. Each observation $z_{target_{i}}\in L\cup\text{Null}$ - is a detection of the object $i$’s location or Null based on the detector’s output for object $i$.
@@
    $=\prod_{i=1}^{n}\Pr(z_{target_{i}}|s_{target_{1}},\ldots,s_{target_{n}},s_{\text{r}})\quad\text{(all }z_{target_{i}}\text{ are independent)}$ (3)
[…]
@@
    • $\textit{Move}_{AB}$ : The cost of moving from location A to B $Cost=-1*N_{a}$ (where $N_{a}$ number of required actions)].